Circular No. 6844

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998V IN NGC 6627 S. Jha, P. Garnavich, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum obtained at the 1.5-m Tillinghast telescope at Mount Hopkins on Mar. 19.5 UT by D. Koranyi and M. Calkins show that SN 1998V (cf. IAUC 6841) is a type-Ia supernova at maximum. The strong Si II absorption shows a photospheric expansion velocity of −10 800 km/s. Narrow emission lines from the host galaxy give a redshift of 5280 km/s.

SUPERNOVA 1998W IN NGC 3075 Jha et al. also report that a spectrum obtained at the 1.5-m Tillinghast

telescope on Mar. 19.4 UT shows that this object (cf. IAUC 6843) is a type-II supernova. A broad Hα feature with a P-Cyg profile is clearly seen despite the low-signal-to-noise data. The photospheric velocity estimated from the minimum of the absorption is −7200 km/s. Narrow emission lines from the host galaxy give a redshift of 3490 km/s.

SUPERNOVA 1998T IN IC 694 D. Balam, University of Victoria, reports the following precise position

for SN 1998T, obtained with the 1.82-m Plaskett telescope of the National Research Council of Canada: α = 11^{h}28^{m}35^{s}.97, δ = +58^{o}33′35′′.9 (equinox 2000.0). A nearby star (R = 17.3) has position end figures 34^{s}.00, 31′21′′.4.

Visual magnitude estimates: Mar. 15.795 UT, 14.3 (K. Hornoch, Lelekovice, Czech Republic); 19.772, 14.3 (Hornoch); 19.776, 14.4 (M. Plsek, Lelekovice, Czech Republic); 20.96, 13.9 (K. Sárneczky, Ágasvár, Hungary).

COMET 29P/SCHWASSMANN-WACHMANN 1 Total visual magnitude estimates: Feb. 20.37 UT, 13.6 (R. Keen, Mt.

Thorodin, CO, 0.32-m reflector); 28.03, 14.3 (J. Carvajal, Avila, Spain, 0.45-m reflector); Mar. 20.44, 12.8 (N. Biver, Honolulu, HI, 0.26-m reflector); 20.96, 12.5 (K. Sárneczky, Ágasvár, Hungary, 0.44-m reflector); 21.05, 12.8 (Carvajal, 0.32-m reflector); 21.06, 12.6 (J. M. San Juan, Avila, Spain, 0.32-m reflector).

1998 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green